Find the area of the surface that is generated by revolving the portion of

the curve y = x


3 between x = 0 and x = 1 about the x −axis.


1
Expert's answer
2021-08-03T13:07:24-0400
"V=\pi\displaystyle\int_{0}^{1}(x^3)^2dx=\pi\big[\dfrac{x^7}{7}\big]\begin{matrix}
 1 \\
 0
\end{matrix}=\dfrac{\pi}{7} (units^3)"

"V=\dfrac{\pi}{7}" cubic units.